OpenStudy (anonymous):

simplify 2a(b^2c)^4 / (2ab^-1)^2

OpenStudy (anonymous):

a. b^10c^4 / 2a b. b^6c^4 / 2a c. b^10c^4 / 4a d. b^10c / 2a

OpenStudy (anonymous):

\[\frac{2a{(b^2c)}^4}{{(2ab^{-1})}^2} = \frac{2ab^8c^4}{4a^2b^{-2}} = \frac{b^8b^2c^4}{2a} = \frac{b^{10}c^4}{2a} \\ \:\\ \textrm{So it's A} \]